I have tried to do this word problem a couple of times, i have a very hard time just understanding where to even begin i draw a complete blank even with my notes, please help.
The area of a square, in square units, with side
x is given by the function A = x2. Find the
area of a square patio that is 40 feet long on a
side.

Answer by jim_thompson5910(35256)

The side is 40 feet long, so the total area of the square is 40^2 = 1600 square feet.

Using the formula, you would get

A = x^2

A = 40^2

A = 1600